OpenStudy (anonymous):

A bin contains a total of 20 batteries, of which 6 are defective. If you select two at random, without replacement, determine the probability that: a. none of the batteries is good. b. at least one battery is good.

OpenStudy (perl):

do you know how to start this

OpenStudy (anonymous):

not exactly

OpenStudy (perl):

lets use combination notation

OpenStudy (perl):

a) in the 20 batteries, there are 6 defective and 14 functional. none of them are good. so you know that both are picked from the defective set

OpenStudy (perl):

( 6 choose 2 ) * (14 choose 0) / ( 20 choose 2 )

OpenStudy (perl):

14 choose zero since we are not picking any from the set of 14 functional

OpenStudy (anonymous):

Ok so (6 C 2) * (14 C 0) / (20 C 2), but we are trying to determine the chance/probability of a. draw a battery the is not functional, and b. at least one battery is functional.

OpenStudy (perl):

right, thats for a)

OpenStudy (perl):

for b) you want probability of choosing 1 defective + probability of choosing 2 defective. it is easier if you use complement

OpenStudy (perl):

it is easier to do the complement

OpenStudy (perl):

let X = # of defective when you choose 2 possibilities for X is X = 0,1,2

OpenStudy (perl):

a) P(X = 0 ) b) P ( X >= 1)

OpenStudy (perl):

now for b) we can do P(X >= 1 )= 1 - P( x < 1) but P( X < 1) = P( X = 0) so b) 1 - P( X = 0) , which we got from part a)